# Probabilistic estimates of the diameters of the Rubik’s Cube groups

### v5

@@ Abstract
The diameter of the Cayley graph of the Rubik’s Cube group is the fewest number of turns needed to solve the Cube from the hardest initial configuration. For the 2$\times$2$\times$2 Cube, the diameter is 11 in the half-turn metric, 14 in the quarter-turn metric, 19 in the semi-quarter-turn metric, and 10 in the bi-quarter-turn metric. For the 3$\times$3$\times$3 Cube, the diameter was determined by Rokicki et al.  to be 20 in the half-turn metric and 26 in the quarter-turn metric. This study shows that a modified version of the coupon collector’s problem in probabilistic theory can predict the diameters correctly for both 2$\times$2$\times$2 and 3$\times$3$\times$3 Cubes insofar as the quarter-turn metric is adopted. In the half-turn metric, the diameters are overestimated by one and two, respectively, for the 2$\times$2$\times$2 and 3$\times$3$\times$3 Cubes, whereas for the 2$\times$2$\times$2 Cube in the semi-quarter-turn and bi-quarter-turn metrics, they are overestimated by two and underestimated by one, respectively. Invoking the same probabilistic logic, the diameters of the 4$\times$4$\times$4 and 5$\times$5$\times$5 Cubes are predicted to be 48 (41) and 68 (58) in the quarter-turn (half-turn) metric, whose precise determinations are far beyond reach of classical supercomputing. The probabilistically estimated diameter is shown to obey the approximate semi-analytical formula of $\ln N/\ln r+\ln N/r$, where $N$ is the number of configurations and $r$ is the branching ratio.
@@ I Objective
The diameter of the Cayley graph of the Rubik’s CubeHofstadter (1981); Singmaster (1981); Biggs (1993); Joyner (2008); Ferenc (2024); West (2019) group (hereafter the diameter) is the fewest number of turns to solve the Cube in the hardest initial configuration. This number is also colloquially referred to as God’s number. Here, a configuration denotes a unique arrangement of sticker colors of the Cube reachable by a series of turns from the solved configuration, while holding the orientation of the whole Cube fixed.
After decades of research by the community of mathematicians and computer scientists,Thistlethwaite (1981); Kociemba (1992); Kunkle and Cooperman (2007, 2009) culminating in a Herculean computational effort by Rokicki et al.,Rokicki et al. (2013, 2014a, 2014b) the diameter of the 3$\times$3$\times$3 Cube was determined to be 20 in the half-turn metric or 26 in the quarter-turn metric.Rokicki et al. (2014b, 2013, a) The corresponding number for the 2$\times$2$\times$2 Cube is 11 in the half-turn metric and 14 in the quarter-turn metric.Ferenc (2024) See below for the definitions of different metrics and other technical details.
@@ II Background
Rubik’s CubeHofstadter (1981); Singmaster (1981); Joyner (2008); Ferenc (2024) has long been a popular testbed for computer science research since the pioneering study by Korf,Korf (1983, 1985) especially in the areas of artificial intelligence and optimization theory. It has also fascinated mathematicians working on group theoryZassenhaus (1982); Fiat et al. (1989); Joyner (2008); Demaine et al. (2011); Ryabogin (2012); Bonzio et al. (2017, 2018) and graph theory.Beenker and Vanlint (1988); Biggs (1993); Ferrero and Hanusch (2014); West (2019) Furthermore, its similarities with problems appearing in quantum mechanics,Corli et al. (2021) statistical mechanics,Lee and Huang (2008); Chen and Ding (2012); Chen and Lee (2014) high-energy physics,Czech et al. (2011) and cryptographyDiaconu and Loukhaoukha (2013); Ionesco and Diaconu (2015) have been noted.
@@
The second group’s objective is to discover or develop an optimization method that is powerful enough to solve the Cube in any configurations efficiently. The method may also be general enough to do so without any domain knowledge, so that it can be applied to other puzzles. Machine Learning algorithms tend to failLichodzijewski and Heywood (2011); Smith et al. (2016); Johnson (2018) because a required fitness function is hard to identify or define. Successful strategies first discover a good fitness function computationallyJohnson (2021) and/or to rely on large pattern databases,Korf (1997); Culberson and Schaeffer (1998); Korf and Felner (2002); Arfaee et al. (2011) both of which are not meant to help a human player. So far the most effective method—the reinforcement learning algorithm recently developed by Agostinelli et al.Agostinelli et al. (2019)—can solve the 3$\times$3$\times$3 Cube 100% of cases; with 60% of cases it discovers the most efficient solutions, underscoring its potential utility in determining the diameters of higher-order Cubes.
The third group computationally searches for the so-called “macros” or a series of turns that a human player can potentially master in a stepwise solution of the Cube. Such a solution is neither expected nor required to be efficient, but it may possibly contain some general principles or strategies for solutions, which may be applied to higher-order Cubes or other related puzzles. These studies, therefore, belong to the developments of expert systemsJackson (1998) or discovery systems.Shen (1990); Gil et al. (2014) Korf’s groundbreaking computational studyKorf (1985, 1983) on Rubik’s Cubes falls into this category, which exhaustively and systematically lists macros for both 2$\times$2$\times$2 and 3$\times$3$\times$3 Cubes. As a by-product, they also set the (albeit conservative) upper bounds for the diameters, spurring many subsequent studies.
@@ III Probabilistic estimate
We shall instead estimate this number by simulating the above brute-force algorithm under the assumption that all generated configurations are random and independent of the previous “seed” configurations (this non-Markovian assumption is clearly not obeyed in reality and is a major approximation). We can then view this problem as a modified version of the coupon collector’s problemGumbel (1941); Dawkins (1991); Flajolet et al. (1992); Blom et al. (1994) in probabilistic theory (see Appendix A). See pages 26 and 34 of Singmaster’s book,Singmaster (1981) where the author develops a similar argument. It is also related to mathematical models of epidemicsHill (2020) and network theory.Newman (2008)
The coupon collector’s problem asks the following question: Let $T_{N}$ be the number of uniformly random integers in the range of 1 through $N$ that need to be generated until every integer 1 through $N$ is realized at least once. What is the expectation value of $T_{N}$? The answer is
@@
The central limit theorem does not apply to $T_{N}$, whose distribution is said to be the Siobhan distribution,Dawkins (1991) which is related to the Gumbel distributions.Gumbel (1941) As per Dawkins,Dawkins (1991) the probability of all integers occurring by the time $T$ random integers have been generated is given asymptotically as
@@
In the $t$th step, from each of the $S(t-1)$ non-redundant seed configurations, $r$ new configurations are generated by performing all $k$ turns, but excluding some combinations of turns (“spurious turns”; see the following sections for examples) that predictably and immediately give rise to duplicate configurations:
@@
Collecting only the non-redundant configurations among the $C(t)$ generated ones, we obtain a new set of seed configurations used in the next step. The size $S(t)$ of this non-redundant set is related to the size $C(t)$ of the original set by
@@
Because the seed set is non-redundant, its size $S(t)$ cannot exceed $N$. This is ensured by Eq. (5). Furthermore, in earlier steps, $C(t)\ll N$ and because the use of $r$ effectively eliminates spurious turns, there is a vanishing probability that two generated configurations are the same, implying $S(t)\approx C(t)$. This is also consistent with the first-order Taylor-series approximation of the exponential in Eq. (5).
@@
will be compared with $E[T_{N}]$. When $T(t)$ exceeds $E[T_{N}]$ for the first time, $t$ is the probabilistic best estimate of the diameter. Nowhere in this estimation do actual configurations need to be generated; nor is a random number needed. The whole analysis proceeds by computing only the estimated sizes of various sets of configurations.
@@ IV.1 Half-turn metric
With the front-top-left cubie held fixed, only the right ($R$), downward ($D$) and back ($B$) layers can be rotated. In the half-turn metric, the Cube can undergo one of the three (3) clockwise 90 turns ($R$, $D$, $B$), three (3) counterclockwise 90 turns ($R^{-1}$, $D^{-1}$, $B^{-1}$), and three (3) 180 turns ($R^{2}$, $D^{2}$, $B^{2}$) in the Singmaster notation.Singmaster (1981)
@@
Table 1: The probabilistic estimates of the numbers of non-redundant seed configurations $S(t)$, redundant configurations $C(t)$, and cumulative redundant configurations $T(t)$ as a function of step ($t$) in the half-turn metric of the 2$\times$2$\times$2 Cube. $N=3674160$, $E[T_{N}]=15.7\,N$, and therefore the predicted diameter is 12 for $T(12)>E[T_{N}]>T(11)$, while the correct diameter is 11.
@@
Table 1 lists $S(t)$, $C(t)$, and $T(t)$ thus computed as a function of $t$ up to $t=12$, where the cumulative number of configurations $T(12)=20.8\,N$ exceeds $E[T_{N}]=15.7\,N$ for the first time. Therefore, the predicted diameter of the 2$\times$2$\times$2 Cube in the half-turn metric is 12, which overshoots the correct diameter of 11.Ferenc (2024)
@@ IV.2 Quarter-turn metric
suggesting $r=120/27=4.44$. Using this $r$ in conjunction with Eqs. (4)–(6), we obtain Table 2. $T(t)$ exceeds $E[T_{N}]=15.7\,N$ for the first time at $t=14$, and therefore the predicted diameter of the 2$\times$2$\times$2 Cube in the quarter-turn metric is 14, which agrees with the correct diameter of 14.Ferenc (2024)
@@ IV.4 Bi-quarter-turn metric
In this another new metric, all nine (9) turns in the half-turn metric plus six (6) additional bi-quarter turns ($RD$, $R^{-1}D^{-1}$, $DB$, $D^{-1}B^{-1}$, $BR$, $B^{-1}R^{-1}$) are elementary operations. This metric is intentionally highly redundant. Lacking some inverses, its “group” and “Cayley graph” may not meet the proper mathematical definitions, but the “diameter” is still well defined.
@@ V The 3$\times$3$\times$3 Cube
The first fraction denotes the number of ways in which the eight corner cubies can be positioned ($8!$) and oriented ($3^{8}$) divided by the three orbits,Joyner (2008) only one of which maintains the sanity.Hofstadter (1981) The second fraction is the number of ways in which the twelve edge cubies can be placed ($12!$) and oriented ($2^{12}$) divided by the four orbits,Joyner (2008) only one of which has the sanity.Hofstadter (1981) The expected number of random configurations to generate all $N$ configurations is therefore
@@ V.1 Half-turn metric
In the first step of the half-turn metric, the 3$\times$3$\times$3 Cube can undergo any one of the following eighteen (18) turns: clockwise 90 face layer turns $R$, $D$, $B$, left ($L$), upward ($U$), front ($F$), and their inverses $R^{-1}$, $D^{-1}$, $B^{-1}$, $L^{-1}$, $U^{-1}$, $F^{-1}$ as well as 180 turns $R^{2}$, $D^{2}$, $B^{2}$, $L^{2}$, $U^{2}$, $F^{2}$ in the Singmaster notation.Singmaster (1981) In the second and subsequent steps, $X$, $X^{-1}$, and $X^{2}$ immediately following the same (where $X=R$, $D$, $B$, $L$, $U$, or $F$) only produce the configurations that are already visited. Furthermore, repeated commutative turns such as $RL$ and $LR$ lead to the same configurations, together indicating $r<15$.
With explicit evaluation,Rokicki et al. (2013) we find
@@
suggesting $r=3240/243=13.33$. Table 5 lists $S(t)$, $C(t)$, and $T(t)$ as a function of $t$. For small $t$, $S(t)$ and $C(t)$ coincide with the “positions” listed in Table 5.1 of Ref. Rokicki et al., 2013, but they are distinct quantities and deviate from each other as $t$ increases. At $t=22$, $T(t)$ exceeds $E[T_{N}]$ for the first time and by a wide margin and, therefore, 22 is the predicted value of the diameter, which overestimates the correct value of 20 (Ref. Rokicki et al., 2013) by two.
@@ V.2 Quarter-turn metric
suggesting $r=1068/114=9.37$. Table 6 predicts how these numbers grow in subsequent steps. It indicates that at $t=26$, $T(t)$ is expected to exceed $E[T_{N}]$ for the first time and hence the predicted diameter is 26, which agrees with the correct diameter determined by Rokicki et al.Rokicki et al. (2013)
@@
Therefore, for the quarter-turn metric, just as in the case of the 2$\times$2$\times$2 Cube, the probabilistic estimate proves to be exact (see more on this in Sec. VIII.1).

### v6

@@ Abstract
The diameter of the Cayley graph of the Rubik’s Cube group is the fewest number of turns needed to solve the Cube from the hardest initial configuration. For the 2$\times$2$\times$2 Cube, the diameter is 11 in the half-turn metric, 14 in the quarter-turn metric, 19 in the semi-quarter-turn metric, and 10 in the bi-quarter-turn metric. For the 3$\times$3$\times$3 Cube, the diameter was determined by Rokicki et al.  to be 20 in the half-turn metric and 26 in the quarter-turn metric. This study shows that a modified version of the coupon collector’s problem in probabilistic theory can predict the diameters correctly for both 2$\times$2$\times$2 and 3$\times$3$\times$3 Cubes insofar as the quarter-turn metric is adopted. In the half-turn metric, the diameters are overestimated by one and two, respectively, for the 2$\times$2$\times$2 and 3$\times$3$\times$3 Cubes, whereas for the 2$\times$2$\times$2 Cube in the semi-quarter-turn and bi-quarter-turn metrics, they are overestimated by two and underestimated by one, respectively. Invoking the same probabilistic logic, the diameters of the 4$\times$4$\times$4 and 5$\times$5$\times$5 Cubes are predicted to be 48 (41) and 68 (58) in the quarter-turn (half-turn) metric, whose precise determinations are far beyond reach of classical supercomputing. The probabilistically estimated diameter is shown to obey the approximate formula of $\ln N/\ln r+\ln N/r$, where $N$ is the number of configurations and $r$ is the branching ratio.
@@ I Objective
The diameter of the Cayley graph of the Rubik’s Cube group Hofstadter (1981); Singmaster (1981); Joyner (2008); Ferenc (2024) (hereafter the diameter) is the fewest number of turns to solve the Cube from the hardest initial configuration. This number is also colloquially referred to as God’s number. Here, a configuration denotes a unique arrangement of sticker colors of the Cube reachable by a series of turns from the solved configuration, while holding the orientation of the whole Cube fixed.
After decades of research by the community of mathematicians and computer scientists Thistlethwaite (2024); Kociemba (2024); Kunkle and Cooperman (2007, 2009), culminating in a Herculean computational effort by Rokicki et al. Rokicki et al. (2013, 2014, 2024), the diameter of the 3$\times$3$\times$3 Cube was determined to be 20 in the half-turn metric or 26 in the quarter-turn metric Rokicki et al. (2024, 2013, 2014). The corresponding number for the 2$\times$2$\times$2 Cube is 11 in the half-turn metric and 14 in the quarter-turn metric Ferenc (2024). See below for the definitions of different metrics and other technical details.
@@ II Background
Rubik’s Cube Hofstadter (1981); Singmaster (1981); Joyner (2008); Ferenc (2024) has long been a popular testbed for computer science research since the pioneering study by Korf Korf (1983, 1985), especially in the areas of artificial intelligence and optimization theory. It has also fascinated mathematicians working on group theory Zassenhaus (1982); Fiat et al. (1989); Joyner (2008); Demaine et al. (2011); Ryabogin (2012); Bonzio et al. (2017, 2018) and graph theory Beenker and Vanlint (1988); Bollobás (2001); Biggs (1993); Ferrero and Hanusch (2014); West (2019). Furthermore, its similarities with problems appearing in quantum mechanics Corli et al. (2021), statistical mechanics Lee and Huang (2008); Chen and Ding (2012); Chen and Lee (2014), high-energy physics Czech et al. (2011), and cryptography Diaconu and Loukhaoukha (2013); Ionesco and Diaconu (2015) have been noted.
@@
The second group’s objective is to discover or develop an optimization method that is powerful enough to solve the Cube in any configurations efficiently. The method may also be general enough to do so without any domain knowledge, so that it can be applied to other puzzles. Machine Learning algorithms tend to fail Lichodzijewski and Heywood (2011); Smith et al. (2016); Johnson (2018) because a required fitness function is hard to identify or define. Successful strategies first discover a good fitness function computationally Johnson (2021) and/or to rely on large pattern databases Korf (1997); Culberson and Schaeffer (1998); Korf and Felner (2002); Arfaee et al. (2011), both of which are not meant to help a human player. So far the most effective method—the reinforcement learning algorithm recently developed by Agostinelli et al. Agostinelli et al. (2019)—can solve the 3$\times$3$\times$3 Cube 100% of cases; with 60% of cases it discovers the most efficient solutions, underscoring its potential utility in determining the diameters of higher-order Cubes.
The third group computationally searches for the so-called “macros” or a series of turns that a human player can potentially master in a stepwise solution of the Cube. Such a solution is neither expected nor required to be efficient, but it may possibly contain some general principles or strategies for solutions, which may be applied to higher-order Cubes or other related puzzles. These studies, therefore, belong to the developments of expert systems Jackson (1998) or discovery systems Shen (1990); Gil et al. (2014). Korf’s groundbreaking study Korf (1985, 1983) falls into this category, which exhaustively and systematically lists macros for both 2$\times$2$\times$2 and 3$\times$3$\times$3 Cubes. As a by-product, they also set the (albeit conservative) upper bounds for the diameters, spurring many subsequent studies.
@@ III Probabilistic estimate
We shall instead estimate this number by simulating the above brute-force algorithm under the assumption that all generated configurations are random and independent of the previous “seed” configurations (this non-Markovian assumption is clearly not obeyed in reality and is a major approximation). We can then view this problem as a modified version of the coupon collector’s problem Gumbel (1941); Dawkins (1991); Flajolet et al. (1992); Blom et al. (1994) in probabilistic theory (see Appendix A). See pages 26 and 34 of Singmaster’s book Singmaster (1981), where the author develops a similar argument. It is also related to mathematical models of epidemics Hill (2020) and network theory Newman (2008).
The coupon collector’s problem asks the following question: Let $T_{N}$ be the number of times an integer in the range of 1 through $N$ is generated uniformly randomly until every integer 1 through $N$ is realized at least once. What is the expectation value of $T_{N}$? The answer is
@@
The central limit theorem does not apply to $T_{N}$, whose distribution is said to be the Siobhan distribution Dawkins (1991), which is related to the Gumbel distributions Gumbel (1941). As per Dawkins Dawkins (1991), the probability that all integers are generated after random generations have been repeated $T$ times is given asymptotically by
@@
In the $t$th step, from each of the $S(t-1)$ non-redundant seed configurations, $r$ new configurations are generated by performing all $k$ turns, but excluding some combinations of turns (“spurious turns”; see the following sections for examples) that immediately give rise to duplicate configurations:
@@
Collecting only the distinct configurations among the $C(t)$ generated ones, we obtain a new set of seed configurations used in the next step. The size $S(t)$ of this non-redundant set is related to the size $C(t)$ of the original set by
@@
Because the seed set is non-redundant, its size $S(t)$ cannot exceed $N$. This is ensured by Eq. (5). Furthermore, in earlier steps, $C(t)\ll N$ and because the use of $r$ effectively eliminates spurious turns, there is a vanishing probability that two randomly generated configurations happen to be the same, implying $S(t)\approx C(t)$. This is also consistent with the first-order Taylor-series approximation of the exponential in Eq. (5).
@@
will be compared with $E[T_{N}]$. When $T(t)$ exceeds $E[T_{N}]$ for the first time, $t$ is the probabilistic best estimate of the diameter. Nowhere in this estimation do actual configurations need to be generated; nor is a random number needed. The whole analysis can in principle proceed by computing only the estimated sizes of various sets of configurations.
@@ IV.1 Half-turn metric
With the front-top-left cubie held fixed, only the right ($R$), downward ($D$) and back ($B$) layers can be rotated. In the half-turn metric, the Cube can undergo one of the three (3) clockwise 90 turns ($R$, $D$, $B$), three (3) counterclockwise 90 turns ($R^{-1}$, $D^{-1}$, $B^{-1}$), and three (3) 180 turns ($R^{2}$, $D^{2}$, $B^{2}$) in the Singmaster notation Singmaster (1981).
@@
Table 1: The probabilistic estimates of the numbers of non-redundant seed configurations $S(t)$, redundant configurations $C(t)$, and cumulative redundant configurations $T(t)$ as a function of step ($t$) in the half-turn metric of the 2$\times$2$\times$2 Cube. $N=3.67\times 10^{6}$, $E[T_{N}]=15.7\,N$, and therefore the predicted diameter is 12 for $T(12)>E[T_{N}]>T(11)$, while the correct diameter is 11.
@@
Table 1 lists $S(t)$, $C(t)$, and $T(t)$ thus computed as a function of $t$ up to $t=12$, where the cumulative number of configurations $T(12)=20.8\,N$ exceeds $E[T_{N}]=15.7\,N$ for the first time. Therefore, the predicted diameter of the 2$\times$2$\times$2 Cube in the half-turn metric is 12, which overshoots the correct diameter of 11 Ferenc (2024).
@@ IV.2 Quarter-turn metric
suggesting $r=120/27=4.44$. Using this $r$ in conjunction with Eqs. (4)–(6), we obtain Table 2. $T(t)$ exceeds $E[T_{N}]=15.7\,N$ for the first time at $t=14$, and therefore the predicted diameter of the 2$\times$2$\times$2 Cube in the quarter-turn metric is 14, which agrees with the correct diameter of 14 Ferenc (2024).
@@ IV.4 Bi-quarter-turn metric
In this another new metric, all nine (9) turns in the half-turn metric plus six (6) additional bi-quarter turns ($RD$, $R^{-1}D^{-1}$, $DB$, $D^{-1}B^{-1}$, $BR$, $B^{-1}R^{-1}$) are elementary operations. This metric is intentionally highly redundant. Lacking some inverses, its “group” and “Cayley graph” may not meet the proper mathematical definitions, but we still use the “diameter” to refer to the fewest number of turns to solve the Cube.
@@ V The 3$\times$3$\times$3 Cube
The first fraction denotes the number of ways in which the eight corner cubies can be positioned ($8!$) and oriented ($3^{8}$) divided by the three orbits Joyner (2008), only one of which maintains the sanity Hofstadter (1981). The second fraction is the number of ways in which the twelve edge cubies can be placed ($12!$) and oriented ($2^{12}$) divided by the four orbits Joyner (2008), only one of which has the sanity Hofstadter (1981). The expected number of random configurations to generate all $N$ configurations is therefore
@@ V.1 Half-turn metric
In the first step of the half-turn metric, the 3$\times$3$\times$3 Cube can undergo any one of the following eighteen (18) turns: clockwise 90 face layer turns $R$, $D$, $B$, left ($L$), upward ($U$), front ($F$), and their inverses $R^{-1}$, $D^{-1}$, $B^{-1}$, $L^{-1}$, $U^{-1}$, $F^{-1}$ as well as 180 turns $R^{2}$, $D^{2}$, $B^{2}$, $L^{2}$, $U^{2}$, $F^{2}$ in the Singmaster notation Singmaster (1981). In the second and subsequent steps, $X$, $X^{-1}$, and $X^{2}$ immediately following the same (where $X=R$, $D$, $B$, $L$, $U$, or $F$) only produce the configurations that are already visited. Furthermore, repeated commutative turns such as $RL$ and $LR$ lead to the same configurations, together indicating $r<15$.
With explicit evaluation Rokicki et al. (2013), we find
@@
suggesting $r=3240/243=13.33$. Table 5 lists $S(t)$, $C(t)$, and $T(t)$ as a function of $t$. For small $t$, $S(t)$ and $C(t)$ coincide with the “positions” listed in Table 5.1 of Ref. Rokicki et al. (2013), but they are distinct quantities and deviate from each other as $t$ increases. At $t=22$, $T(t)$ exceeds $E[T_{N}]$ for the first time and by a wide margin and, therefore, 22 is the predicted value of the diameter, which overestimates the correct value of 20 (Ref. Rokicki et al. (2013)) by two.
@@ V.2 Quarter-turn metric
suggesting $r=1068/114=9.37$. Table 6 predicts how these numbers grow in subsequent steps. It indicates that at $t=26$, $T(t)$ is expected to exceed $E[T_{N}]$ for the first time and hence the predicted diameter is 26, which agrees with the correct diameter determined by Rokicki et al. Rokicki et al. (2013).
@@
Therefore, for the quarter-turn metric, just as in the case of the 2$\times$2$\times$2 Cube, the probabilistic estimate proves to be exact.
V.3 Square-turn metric
The square-turn metric Singmaster (1981); Joyner (2008) allows six (6) $180^{\circ}$ turns $R^{2}$, $D^{2}$, $B^{2}$, $L^{2}$, $U^{2}$, and $F^{2}$. The number of configurations in this metric is much fewer Joyner (2008).
$N={4!4}\frac{4!4!4!}{2}=6.63\times 10^{5}.$ (64)
The first factor of $4!$ is the number of ways in which the urf corner cubie Singmaster (1981) and its three diagonal ones (ulf, drb, dlb) can interchange their positions by the $R^{2}$, $D^{2}$, and $B^{2}$ turns (their orientations are fixed by their positions). The second factor of $4$ is the number of positions the ulf cubie can reach by the $L^{2}$, $U^{2}$, and $F^{2}$ turns. The remaining three corner cubies’ positions and orientations are uniquely determined by the first five. The quotient $4!4!4!/2$ is the number of ways three groups of 4 edge cubies (sharing a common center slice) can be positioned with their orientations fixed. Since the number of correctly oriented edge cubies is always even, there are two orbits and the number must be divided by two.
The expected number of random configurations to be generated to reach the diameter is
$E[T_{N}]\pm\sigma[T_{N}]=\Big{(}14.0+1.3\Big{)}\,N.$ (65)
This group has a striking similarity with the 2$\times$2$\times$2 Cube group in the quarter-turn metric (Sec. IV.2) with the only difference being the number of configurations. Their $S(t)$ and $C(t)$ therefore display the identical behavior for lower $t$.
$S(0)=C(0)$ $=$ $1,$ (66)
$S(1)=C(1)$ $=$ $6,$ (67)
$S(2)=C(2)$ $=$ $27,$ (68)
$S(3)=C(3)$ $=$ $120$ (69)
The branching ratio ($r=120/27=4.44$) is also the same.
Table 7: The probabilistic estimates of the numbers of non-redundant seed configurations $S(t)$, redundant configurations $C(t)$, and cumulative redundant configurations $T(t)$ as a function of step ($t$) in the square-turn metric of the 3$\times$3$\times$3 Cube. $N=6.63\times 10^{5}$, $E[T_{N}]=14.0\,N$. The predicted diameter is 13 for $T(13)>E[T_{N}]>T(12)$, while the correct diameter is 15.
$t$ $S(t)$ $C(t)$ $T(t)$
0 1 1 1
1 6 6 7
2 27 27 34
3 120 120 154
4 0.0008 $N$ 0.0008 $N$ 0.0010 $N$
5 0.0036 $N$ 0.0036 $N$ 0.0046 $N$
6 0.0157 $N$ 0.0158 $N$ 0.0204 $N$
7 0.0672 $N$ 0.0696 $N$ 0.0900 $N$
8 0.2580 $N$ 0.2984 $N$ 0.3884 $N$
9 0.6820 $N$ 1.1456 $N$ 1.5340 $N$
10 0.9516 $N$ 3.0279 $N$ 4.5618 $N$
11 0.9854 $N$ 4.2250 $N$ 8.7869 $N$
12 0.9874 $N$ 4.3751 $N$ 13.1619 $N$
13 0.9875 $N$ 4.3841 $N$ 17.5460 $N$
14 0.9875 $N$ 4.3846 $N$ 21.9307 $N$
15 0.9875 $N$ 4.3846 $N$ 26.3153 $N$
Table 7 estimates the subsequent growth and decay of these numbers. It can be seen that $T(t)$ exceeds $E[T_{N}]$ for the first time at $t=13$. This predicted diameter (13) is smaller by as much as two than the correct diameter of 15, which has been determined by brute-force computation. Worse yet, as per Eq. (3), the probability that the predicted diameter is actually 12 instead of 13 is not negligible.
$\text{Pr}(T_{N}\leq T(12))$ $=$ $0.279,$ (70)
$\text{Pr}(T_{N}\leq T(13))$ $=$ $0.984$ (71)
In other words, there is a fair chance that one last configuration is yet to be realized after 12 steps.
$Ne^{-T(12)/N}$ $=$ $1.3,$ (72)
$Ne^{-T(13)/N}$ $=$ $0.016,$ (73)
After 13 steps, all configurations should have been generated according to our estimates. Hence, this metric turns out to be the hardest case for the probabilistic estimates.